7. Calculate the pH of 3.00 x 10-4 M soln of codeine if its pK, of the conjugate acid is 8.00. Answer: pH =
Answer #1

Feel free to ask doubt and kindly rate it

For de acid, 10g 3=0.6 any weak pka=8 (giren) I c = concentration pH = pka-loge 2 - log [3x10-9] = 8 - log [ 3x10 -4 ] ] = lg